# Perfect Superconducting Diode and Supercurrent Range Controller

### v1

@@ Abstract
If VSRs overlap, then the device can potentially switch its vorticity state which might be undesirable in some applications. To eliminate the overlap the critical phases has to be as small as possible, although, physically, it is never lower than $\pi/2$. Disjoint VSRs have an advantage in that the device is not able to transition to a different vorticity state without going through the normal state. Since the creation of the normal state requires a significant amount of energy, specifically, the condensation energy, such transitions are exponentially suppressed at low temperaturesSang L. Chu and Bezryadin (2004). By choosing a low value for $\phi_{c}$ one can reduces or even eliminate the VSRs overlap because the area of the VSR gets smaller if $\phi_{c}$ is reduced, as demonstrated below.
@@
If b-invariant PSD is desirable then all the critical phases must be integer multiples of $2\pi/n$, but SRC does not require this condition. For a 3-SQUID b-invariant PSD, if we choose $\phi_{c}=2\pi/3$ such that $(3\phi_{c})/(2\pi)=1$, we obtain disjoint VSRs, i.e, for any choice ($j$, $b$) there is only one possible vorticity sate (Fig.4a). Vorticity states, for which the 2nd wire is the switching wire and which satisfy Eq.7 are $[1,0]$, $[2,1]$, $[3,2]$, etc. For a 2-SQUID SRC, if we choose $\phi_{c,1}=\pi/2$ and $\phi_{c,2}=2.5$ (rad), the resulting VSRs is also disjoint (Fig.4b), i.e., there is always a non-superconducting region between neighbor VSRs.
@@
In conclusion, a model of superconducting quantum interference device with multiple weak links (nanowires) with a linear current-phase relationship is analyzed. The model predicts that such systems can act as perfect superconducting diodes (PSDs) and/or as superconducting range controllers (SRCs). In PSD, The critical current of one polarity is exactly zero and the critical current of the opposite polarity is above zero ($\eta=1$). In SRC, the supercurrent is bounded between two positive values ($\eta>1$). A point-wise PSD device may be realized with 2-SQUIDs whereas a b-invariant PSD device is realized with SQUIDs with $>2$ wires. In contrast, a SRC device may be realized with 2-SQUID. A small difference between the nanowires does not eliminate the strong diode effect reported here, but it changes PSD into SRC.
@@ Acknowledgements.
We acknowledge support from NSF DMR-2104757.

### v3

@@ VI Discussion
If VSRs overlap, then the device can potentially switch its vorticity state which might be undesirable in some applications. To eliminate the overlap the critical phases has to be as small as possible, although, physically, it is never lower than $\pi/2$. Disjoint VSRs have an advantage in that the device is not able to transition to a different vorticity state without going through the normal state. Since the creation of the normal state requires a significant amount of energy, specifically, the condensation energy, such transitions are exponentially suppressed at low temperaturesChu et al. (2004). By choosing a low value for $\phi_{c}$ one can reduces or even eliminate the VSRs overlap because the area of the VSR gets smaller if $\phi_{c}$ is reduced, as demonstrated below.
@@
If b-invariant PSD is desirable then all the critical phases must be integer multiples of $2\pi/n$, but SRC does not require this condition. For a 3-SQUID b-invariant PSD, if we choose $\phi_{c}=2\pi/3$ such that $(3\phi_{c})/(2\pi)=1$, we obtain disjoint VSRs, i.e, for any choice ($j$, $b$) there is only one possible vorticity state (Fig.6a). Vorticity states, for which the 2nd wire is the switching wire and which satisfy Eq.7 are $[1,0]$, $[2,1]$, $[3,2]$, etc. For a 2-SQUID SRC, if we choose $\phi_{c,1}=\pi/2$ and $\phi_{c,2}=2.5$ (rad), the resulting VSRs is also disjoint (Fig.6b), i.e., there is always a non-superconducting region between neighbor VSRs.
@@
VII Conclusion
A model of superconducting quantum interference device with multiple weak links (nanowires) with a linear current-phase relationship is analyzed. The model predicts that such systems can act as perfect superconducting diodes (PSDs) and/or as superconducting range controllers (SRCs). In PSD, The critical current of one polarity is exactly zero and the critical current of the opposite polarity is above zero ($\eta=1$). In SRC, the supercurrent is bounded between two positive values ($\eta>1$). A point-wise PSD device may be realized with 2-SQUIDs whereas a b-invariant PSD device is realized with SQUIDs with >2>2 wires. In contrast, a SRC device may be realized with 2-SQUID. A small difference between the nanowires does not eliminate the strong diode effect reported here, but it changes PSD into SRC.
@@ Acknowledgements.
The work was supported in part by the NSF DMR-2104757 and by the NSF OMA 2016136 Quantum Leap Institute for Hybrid Quantum Architectures and Networks (HQAN). We also like to thank the University of Illinois Materials Research Laboratory for providing us with capabilities of sample fabrication and imaging.